# 1Manifold Research 2Metarch.ai 3Georgia Tech

### v1

@@ Abstract
Vision-language-action (VLA) models represent an important step toward general-purpose robotic systems by integrating visual perception, language understanding, and action execution. However, systematic evaluation of these models, particularly their zero-shot generalization capabilities in procedurally generated, out-of-distribution (OOD) environments, remains limited. In this paper, we introduce MultiNet v0.2, a comprehensive benchmark designed to evaluate and analyze the generalization performance of state-of-the-art VLM and VLA models—including GPT-4o, GPT-4.1, OpenVLA, Pi0 Base, and Pi0 FAST—on diverse procedural tasks from the Procgen benchmark. Our analysis reveals several critical insights: (1) all evaluated models exhibit significant limitations in zero-shot generalization to OOD tasks, with performance heavily influenced by factors such as action representation and task complexity; (2) OpenVLA generally outperforms other models due to its robust architectural design; and (3) GPT variants demonstrate substantial improvements when constrained appropriately, highlighting the sensitivity of model performance to precise prompt engineering.
@@ I Introduction
The architecture and training methodologies of these models play a crucial role in their generalization abilities. Factors such as the nature of the action space (continuous vs. discrete), the domain of training data (real-world robotics vs. simulated environments), and the methods of input-output processing can significantly impact performance [30]. For instance, models trained predominantly on robotics data may not perform well in simulated game environments due to differences in action representation and environmental complexity.
@@
Our contributions are threefold:
  • We provide a systematic benchmarking framework that evaluates the performance of state-of-the-art models on a variety of procedurally generated environments.
  • We analyze the impact of architectural choices ,training data, and output processing techniques on model generalization, highlighting the limitations of current approaches.
  • We offer insights into how factors like action space representation and image complexity influence model performance, paving the way for future improvements in generalist AI systems.
@@ II Related Work
Our work builds on the rapid advancement of physical and digital vision-language-action models that enable agents to interpret, reason, and act across diverse tasks and embodiments. In this section, we review the landscape of related work, organized into four key areas with an emphasis on digital environments: existing action models, datasets and benchmarks, data collection tools, and agentic applications. For comprehensive surveys on VLAs and computer use agents, see [36, 45, 49, 61].
Action models. Vision-language-action (VLA) architectures generally consist of four core components: vision encoders, language encoders, planners, and action decoders. Recent work has advanced each element to improve end-to-end performance.
First, on the encoder and fusion side, richer multimodal representations have been achieved by integrating advanced vision backbones and efficient conditioning mechanisms. For example, SigLIP [82] image embeddings have been adopted in OpenVLA [39] and PaLiGemma [4], while FiLM [55] conditioning enables RT-1 [9], BC-Z [38], and OpenVLA-OFT [40] to modulate visual features with language cues. More recently, Interleave-VLA [21] interleaves vision and language embeddings at multiple layers to improve cross-modal alignment.
Second, the planner, sitting between perception and action, can be categorized into three paradigms.
  • World model planner, which first learns a compact dynamics model in a latent space and then ’imagines’ trajectories for policy extraction. Representative examples include Dreamer series [32], Genie series [20], DINO-WM [85], AdaWorld [25], and Latent Diffusion Planning (LDP) [75].
  • Sequence-based action generation, where control is treated as conditional sequence prediction by large transformers. This category unifies (a) pretrained multimodal decoders, such as RT-2 and OpenVLA, which generate actions autoregressively over vision–language embeddings [10, 39], and diffusion-policy methods like Octo [26], TinyVLA [72], $\pi_{0}$ [6], CogACT [44], and GR00T N1 [5], with (b) pure sequence planners like the Decision Transformer and its hierarchical variants, which autoregressively predict future actions from past state–action histories and a return-to-go token [13, 15].
  • Reactive model-free control, in which policies are learned directly from demonstrations or offline data without any explicit planning module. Classic methods include Behavior Cloning[56], DAgger [60], and GAIL [33] and new variants such as Implicit Behavior Cloning[23], which excel in short-horizon or well-covered tasks but lack the capability for roll-out–based foresight.
Each paradigm balances sample efficiency, generalization capability, and the degree of explicit world modeling: model-based planners offer strong long-horizon reasoning at the cost of model learning effort; sequence-based methods leverage powerful pretrained sequence models to bypass explicit dynamics but depend on large-scale pretraining; and model-free approaches provide simplicity and stability in familiar settings but struggle with unobserved or out-of-distribution scenarios.
Specifically, to handle long-horizon and hierarchical tasks, many VLA systems combine:
  • Speed–accuracy cascades: two-model pipelines with fast/slow predictors[6, 7, 68, 22, 5].
  • Planning precision: chain-of-thought reasoning with VLAs[84].
  • Horizon compression: hierarchical vector quantization in VQ-BeT[42].
  • Action-chunking: parallel decoding in OpenVLA-OFT and FAST token compression[39, 37].
  • Latent planning: AgiBot GO-1 ViLLA framework with a VQ-VAE latent model and a 24-layer transformer[68].
These architectural advances have enabled VLAs to excel not only in standard benchmarks but also across very different application domains. In robotics, OpenVLA [39] is fine-tuned to excel in 29 multi-object manipulation tasks (e.g., block stacking, pick-and-place, table cleaning); VQ-BeT [42] performs multistep kitchen behaviors such as drawer opening/closing and toaster/fridge operation; and the $\pi$-series [6, 37, 7] demonstrate zero-shot generalization on tasks like bowl stacking, towel folding, and grocery bagging. In digital environments, WebGUM [24], OpenAI Operator[54], Claude Computer Use [2], Manus AI [67], and CogAgent [35] can navigate GUIs and carry out web tasks such as clicking, typing, filling in forms, searching, and booking, while Ferret-UI [81] excels at low-level interface understanding (e.g., icon and widget detection). Notably, unified models like Magma [79] that bridge digital GUI actions with physical-world manipulation still remain underexplored.
Datasets & benchmarks. Recent embodied VLA work utilizes large-scale, multitask robot datasets and benchmarks such as OpenX [51], LIBERO [48], BridgeData [18, 19], VLABench [83], Agibot World [68], egocentric benchmarks such as EGO-4D [28], augmented datasets with affordance, action trajectory, and object annotations [16, 79], and video-action datasets such as Something-SomethingV2 [27], Motion-X [46], and MotionBank [77] using actions annotations or relying on architectural components [43, 80] to learn actions from human demonstration videos. In the domain of OS and browser use, GUI agents are evaluated on web pages, desktop UIs with datasets such as Mind2Web (Web navigation tasks) [17], AITW (Android device control) [58], WebUI (400 K rendered pages + metadata) [73], OSWorld (real-computer environment tasks) [76], GUI-World (6 GUI scenarios, dynamic content) [12], and VideoGUI (instructional‐video–driven GUI editing) [47].
Data collection tools. Compared to the abundant text training data on the Internet for natural language models, the lack of action datasets [41] requires proper data collection tools for VLAs. In robotics, some mobile teleoperation interfaces were engineered for tasks such as 6-DoF demonstration [50], and multi-arm coordination [70]. As for computer use agents, commercial RPA tools such as UiPath and Automation 360 capture user interactions with applications, including clicking, reading, and writing actions on user interface objects while open-source browsers (Selenium IDE, Playwright Codegen, Browserless Recorder) capture browser operations. On the OS level, open source tools such as selfspy [63], can capture keystrokes, visited URLs, and executed commands. Even though these emerging tools extend our ability to collect large-scale computer-use action datasets, the tools for action data collection in digital VLAs are still lacking, and building large-scale, high-quality action datasets remains a major challenge.
Agentic applications. Digital agentic systems now operate across desktop and mobile platforms as well as within web browsers, serving domains from customer support, finances, to healthcare and enterprise productivity.
Closed source, generalist multimodal computer use agents such as OpenAI Operator, Claude Computer Use, Manus AI interact with GUIs and browser content using only image and text inputs to automate workflows end-to-end[54, 2, 67]. Open source agents such as AutoGPT [3] and OS-Copilot [74] interface with terminals, file systems, and third-party applications, while specialized web agents [24, 29] focus on browser navigation and form automation.
In enterprise contexts, CRM and support workflows are increasingly driven by agents like Salesforce Einstein GPT [62] and ServiceNow Virtual Agents [65] ground responses in enterprise data to automate case routing, conversation summarization, and triage escalations. In healthcare, tools such as Google’s Articulate Medical Intelligence Explorer (AMIE) [69] demonstrate text-based diagnostic performance on par with or exceeding primary care physicians in controlled studies.
Despite these advances, current LLM-based agentic applications still struggle with reliability, accuracy, and efficiency when faced with dynamic, complex, and uncontrolled real-world environments due to their fundamentally incomplete world model [71], tendency toward self-deception [78], and lack of appropriate bounds on agency [11].
@@ III-A Dataset
For evaluation purposes, we established test splits comprising 10 percent of the episodes randomly sampled from each subdataset. Given that the Procgen dataset consists of 16 distinct subdatasets, this approach resulted in balanced and representative test splits for comprehensive evaluation. Importantly, all profiling and performance assessment experiments reported in this study were conducted exclusively on these designated test splits, ensuring that our results reflect the models’ genuine generalization capabilities without contamination from training data.
@@ III-B Models
OpenVLA: The weights were obtained from the HuggingFace repository under the identifier openvla/openvla-7b.
Pi0 Base: We utilized weights stored at s3://openpi-assets/checkpoints/pi0_base/params. We obtained this path from the Openpi open-source codebase [^3] https://github.com/Physical-Intelligence/openpi
@@ III-B1 Data Ingestion
GPT 4o and GPT 4.1: Employed the Genesis prompt engineering framework [^4] https://github.com/ManifoldRG/MultiNet/tree/main/src/modules, meticulously crafting timestep-specific prompts that including detailed instructions and constraints for inference, explicit descriptions of the subdataset task and environment, visual observations corresponding to the current timestep, clear specifications of the action space and range, and verbal descriptions for each possible action. As seen in the Appendix section VII-C the prompt constructed for GPT 4.1 was improved over the v0.1 Genesis prompt to observe the performance increase of VLMs due to a more rigorously curated prompt. The information given to the VLMs was the same through the old and new prompts, but the new one includes changes to ensure more valid outputs from the model. The new prompt explicitly includes the fact that this is a simulation dataset. The new prompt also uses capitalization to emphasize instructions. Additionally, rather than stating all the possible types of inputs that could be given from different types of action datasets in Multinet, the new prompt includes only the type of inputs that will be given as a part of the Procgen datasets.
@@ III-B2 Model Adaptation
OpenVLA: Adaptation involved restricting the autoregressive step to one, ensuring single-dimensional action vector predictions. Generated actions were unnormalized using statistics computed across the entire Procgen subdataset and subsequently rounded to discrete actions. For Brier MAE calculation, logits from the Llama 2 backbone were extracted, probabilities computed, and token mappings from the Llama vocabulary to Procgen integer actions were created and cached to facilitate efficient inference.
Pi0 Base: The action horizon was constrained to one timestep, with flow matching denoising executed over ten steps (default configuration). Actions were predicted with a default dimension of 32. The first dimension of the predicted action was picked to ensure a final single dimensional action vector prediction, which was then unnormalized based on Procgen dataset-wide statistics (same dataset statistics used for OpenVLA), and discretized by rounding.
Pi0 Fast: Adaptation entailed adjusting the action horizon to one timestep and setting the action dimension to one, compatible with Procgen’s single-dimensional action space. The decoding steps were limited to four tokens (”Action”, ”:”, a space character, and the Paligemma (VLM backbone of Pi0 models) location token) to explicitly generate the correct action value. To compute probabilities for Brier MAE, leveraging to the static nature of the BPE tokenizer, token mappings between the Paligemma token IDs and Procgen integer actions were established post-training. During inference, probabilities extracted from the Paligemma backbone logits were aggregated according to this mapping. Recognizing Pi0 Fast’s inference speed was significantly slower (approximately ten times slower than OpenVLA), we optimized runtime by caching embeddings for the two static zero-images after initial processing by the SigLIP embedding, effectively doubling inference speed.
@@ III-C Inference Infrastructure
  • In contrast, GPT 4x inference was managed externally through batch job APIs, eliminating local processing overhead and leveraging scalable cloud resources to efficiently handle its computational requirements.
@@ III-D Evaluation metrics
In order to capture the ability of state-of-the-art VLAs and VLMs to generalize to completely unseen, OOD action data, we carefully curated a list of metrics to evaluate the models as fairly as possible. These metrics capture a variety of aspects of performance of the models such as how well calibrated they are in their estimates, reliance on outliers, how they get affected by class imbalance, how well they are able to restrict their predictions to the valid action space, whether the models are biased towards certain kinds of subdatasets or actions, etc.
@@
    Brier $\text{Brier score}=\frac{1}{N}\sum_{t=1}^{N}\sum_{i=1}^{R}|f_{ti}-o_{ti}|$ (1)
@@
    Brier MAE is a variation of the original Brier score [8], which is a useful method to measure the accuracy of probabilistic predictions. In our case, the probability that the action in a given timestep is one out of all the possible actions in the action space. It gives a good idea of how well the predictions of the models are calibrated. Brier score penalizes predictions that are both wrong and confident more severely than those that are wrong but uncertain. Additionally, unlike Mean Squared Error, using Mean Absolute Error maintains the original scale of the data making it potentially more interpretable in the action space’s context. MAE is also less sensitive to outlier predictions than squared error metrics, thus giving a more stable evaluation of typical performance. Maximum possible Brier score is 2, which is assigned to invalid predictions as a high penalty.
@@
  • Normalized Quantile Filtered Brier Mean Absolute Error This metric value is obtained by filtering out the Brier Absolute Errors that are greater than or equal to the 5th percentile error or lesser than or equal to the 95th percentile error, and then normalizing based on the quantile filtered minimum and maximum errors. This metric helps indicate how good or bad the majority of the predictions are. Discrepancy in trends between normalized and normalized quantile filtered Brier MAEs indicates the reliance on a few extremely good or extremely bad predictions.
@@
    This metric quantifies how much the worst case error deviates from the typical (median) error produced by the model for a given subdataset of Procgen. A value significantly greater than 1 indicates the presence of some extremely bad predictions that can skew the results.
@@
    For our case, where it is a multi-class, single correct prediction task and the classes (actions) are mutually exclusive, false positives are equal to false negatives. This is because an incorrect predicted action in a given timestep is a false positive for its class and a false negative for the ground truth action class. Due to this, precision = recall = f1 score = exact match rate when calculating micro metrics in our case. The micro precision/recall/f1 is calculated over all timesteps of the test split of a given subdataset. This metric measures how precise the models are - high precision indicates that the model gets a high number of predictions correct, and a low precision suggests the model is frequently predicting actions that don’t match the ground truth.
  • Percentage invalids Predictions that fall outside the valid action space of a given subdataset of Procgen are considered invalids. In the case of the VLMs - GPT 4o, and GPT 4.1, the outputs that are not in the desired format of an action space size vector containing discrete values as its elements, are also considered invalids. Thus, percentage invalids is calculated as
@@
  • Micro Precision without Invalids This metric is a variation of micro precision where the false positives are only counted for valid predictions. By not considering the invalid predictions, this metric gives an idea about the model’s performance in the cases where it is able to produce valid outputs. A discrepancy in the trends between micro precision and micro precision without invalids indicates that a model would be more capable if it was fine-tuned, prompted, or had its output processed to constrain the predictions to the valid action space.
@@
    These metrics help understand model biases and preferences towards specific action classes.
@@
    Macro precision, recall, and F1 metrics are calculated by averaging the class-wise metrics across all classes in the valid action space of a given subdataset for a given model. These metrics are very important as they help alleviate the effects of the dominating majority class that skews performance. With micro metrics, the majority classes are favored due to global aggregation. This leads to a misunderstood notion of performance, especially in the Procgen expert data, as there are numerous cases of class imbalance which can in turn skew the values of the micro metrics. Macro metrics prove to be a better indicator of a model’s performance when compared to micro metrics irrespective of the nature of the test dataset as they treat all classes equally regardless of size. Within this, Macro recall is typically the most representative metric for performance as it is neither affected by majority classes as in the case of micro metrics, nor is it affected by the performance of models on the rare classes as in the case of macro precision or F1 scores.

### v2

# Pranav Guruprasad∗+12, Yangyue Wang∗12, Sudipta Chowdhury∗1, Harshvardhan Sikka123, Paul Pu Liang4

@@ Abstract
Vision-language-action (VLA) models represent an important step toward general-purpose robotic systems by integrating visual perception, language understanding, and action execution. However, systematic evaluation of these models, particularly their zero-shot generalization capabilities in procedurally out-of-distribution (OOD) environments, remains limited. In this paper, we introduce MultiNet v0.2, a comprehensive benchmark designed to evaluate and analyze the generalization performance of state-of-the-art VLMs and VLAs—including GPT-4o, GPT-4.1, OpenVLA, Pi0 Base, and Pi0 FAST—on diverse procedural tasks from the Procgen benchmark. Our analysis reveals several critical insights: (1) all evaluated models exhibit significant limitations in zero-shot generalization to OOD tasks, with performance heavily influenced by factors such as action representation and task complexity; (2) VLAs generally outperforms other models due to their robust architectural design; and (3) VLM variants demonstrate substantial improvements when constrained appropriately, highlighting the sensitivity of model performance to precise prompt engineering. We release our benchmark, evaluation framework, and findings to enable the assessment of future VLA models and identify critical areas for improvement in their application to out-of-distribution digital tasks.
@@ I Introduction
The architecture and training methodologies of these models play a crucial role in their generalization abilities. Factors such as the nature of the action space (continuous vs. discrete), the domain of training data (real-world robotics vs. simulated environments), and the methods of input-output processing can significantly impact performance. For instance, models trained predominantly on robotics data may not perform well in simulated game environments due to differences in action representation and environmental complexity.
@@
Our primary contributions in this paper are:
  • A systematic benchmarking framework that evaluates the performance of state-of-the-art models on a variety of procedurally generated game environments.
  • Detailed profiling results for an initial set of SoTA VLMs and VLAs, along with analysis of their performance.
  • Analysis of the impact of architectural choices, training data, and output processing techniques on model generalization, highlighting the limitations of current approaches.
  • Insights into how factors like action space representation and image complexity influence model performance.
@@
Figure 1: Model rankings across datasets by macro recall, highlighting consistent top performance by OpenVLA and GPT-4.1.
@@ II Related Work
Our work builds on the rapid advancement in vision-language-action (VLA) models for embodied AI, both in physical robots and digital environments, that enable agents to perceive, reason, and act across diverse tasks and embodiments. In this section, we review the landscape of related work, organized into two complementary areas: evaluation in digital environments and evaluation of generalist VLAs. For comprehensive surveys on VLAs and computer use agents, see [16, 19, 21, 28].
Evaluation in digital environments. Digital benchmarks play a crucial role in measuring agents’ decision-making, perception, and planning skills in simulated settings. We categorize prominent digital benchmarks into three groups:
  • Gameplay Benchmarks. Benchmarks such as Gymnasium, which wraps the Arcade Learning Environment (ALE) for Atari games and provides continuous-control tasks, have become standard testbeds for reinforcement-learning agents[2]. More recent procedurally generated game benchmarks such as Procgen, MineDojo and competition and crafting game benchmarks such as AlphaStar, Crafter introduce a family of tasks that probe models’ ability in 2D/3D games with more complex objectives and environments to further assess agents’ ability to adapt and generalize[7, 11, 34, 14].
  • Interface and Web-Interaction Benchmarks. As agent evaluation transitions from controlled gameplay simulations to more general computer use tasks, benchmarks such as Mind2Web, AITW, and OSWorld aim to evaluate proficiency in tasks such as web browsering, menu navigation, and form completion across environments such as web applications, Windows, MacOS, Android, and Linux. [9, 26, 35, 6, 20].
In recent years, agents driven by large language models have been evaluated in subsets of these benchmarks[31, 12].
We extend digital evaluation by systematically including state-of-the-art VLAs alongside VLM-based agents, applying a unified metric suite across all procedurally generated and discrete-control environments.
Evaluation of generalist VLAs. Generalist VLAs are commonly evaluated using in-domain robotics benchmarks such as OpenX, VLABench, and LIBERO, consisting of common physical manipulation tasks and robot embodiments [22, 10, 38, 32, 8]. These assessments focus on metrics such as task completion rates, sim-to-real transfer fidelity, and robustness to sensor noise. Some recent efforts such as Magma [36] have attempted to narrow cross-domain transfer, but none systematically unify digital and physical benchmarks under consistent metrics and datasets.
In MultiNet v0.2, we repurpose robotics-trained VLAs and general-purpose VLMs as digital agents and benchmark them in procedurally generated 2D environments.
@@ III-A Dataset
For evaluation purposes, we established test splits comprising 10% of the episodes randomly sampled from each subdataset. Given that the Procgen dataset consists of 16 distinct subdatasets, this approach resulted in balanced and representative test splits for comprehensive evaluation. Importantly, all profiling and performance assessment experiments reported in this study were conducted exclusively on these designated test splits, ensuring that our results reflect the models’ genuine generalization capabilities without contamination from training data.
@@ III-B Models
OpenVLA: The weights were obtained from the HuggingFace repository under the identifier openvla/openvla-7b. [^3] https://huggingface.co/openvla/openvla-7b
Pi0 Base: We used weights stored at s3://openpi-assets/checkpoints/pi0_base/params. We obtained this path from the Openpi open-source codebase [^4] https://github.com/Physical-Intelligence/openpi
@@ III-B1 Data Ingestion
GPT 4o and GPT 4.1: Employed the Genesis prompt engineering framework, which acts as a crucial intermediary between the data and the model, by translating the raw Procgen trajectory data (images and associated metadata) into a rich, structured textual representation that primes the VLM for analyzable outputs. Further information about the Genesis framework is detailed in Appendix Section VII-C.
@@ III-B2 Model Adaptation
OpenVLA: Adaptation involved restricting the autoregressive step to one, ensuring single-dimensional action vector predictions. Generated actions were unnormalized using statistics computed across the entire Procgen subdataset and subsequently rounded to discrete actions. To compute evaluation metrics such as Brier MAE 1, we first extracted logits from the Llama 2 backbone and computed the corresponding probabilities [33]. We then initialized and cached mappings from Llama vocabulary tokens to Procgen integer actions. During inference, this approach allowed us to group the probabilities of each vocabulary token according to their associated integer action classes after unnormalization, while maintaining efficient inference throughout the process.
Pi0 Base: The action horizon was constrained to one timestep, with flow matching denoising executed over ten steps (default configuration). Actions were predicted with a default dimension of 32. The first dimension of the predicted action was picked to ensure a final single-dimensional action vector prediction, which was then unnormalized based on Procgen dataset-wide statistics (same dataset statistics used for OpenVLA), and discretized by rounding.
Pi0 Fast: Adaptation entailed adjusting the action horizon to one timestep and setting the action dimension to one, compatible with Procgen’s single-dimensional action space. To explicitly generate the correct action value, the decoding steps were limited to four tokens: “Action”, “:”, a space character, and the Paligemma (VLM backbone of Pi0 models) location token. To compute probabilities for Brier MAE, leveraging the static nature of the BPE tokenizer, token mappings between the Paligemma token IDs and Procgen integer actions were established. During inference, probabilities extracted from the Paligemma backbone logits were aggregated according to this mapping. Recognizing Pi0 Fast’s significantly slower inference speed (approximately ten times slower than OpenVLA), we optimized runtime by caching embeddings for the two static zero-images after initial processing by the SigLIP embedding [37], effectively doubling inference speed.
@@ III-C Inference Infrastructure
  • In contrast, GPT 4x inference was managed externally through OpenAI’s Batch API, eliminating local processing overhead and leveraging scalable cloud resources to efficiently handle its computational requirements.
@@ III-D Evaluation metrics
In order to capture the ability of state-of-the-art VLAs and VLMs to generalize to completely unseen, OOD action data, we carefully curated a list of metrics to evaluate the models as fairly as possible. These metrics address diverse aspects of model performance, such as how well calibrated they are in their estimates, reliance on outliers, how they are affected by class imbalance, how well they are able to restrict their predictions to the valid action space, whether the models are biased towards certain kinds of subdatasets or actions, etc.
@@
    Brier $\text{Brier MAE}=\frac{1}{N}\sum_{t=1}^{N}\sum_{i=1}^{R}|f_{ti}-o_{ti}|$ (1)
@@
    Brier MAE is a variation of the original Brier score [4], which is a useful method to measure the accuracy of probabilistic predictions. In our case, for a given timestep and the associated action space, each model computes a probability for each action that represents how likely it is to be equal to the truth label. This measures how well the predictions of the models are calibrated. Brier score penalizes predictions that are both wrong and confident more severely than those that are wrong but uncertain. Additionally, unlike Mean Squared Error, using Mean Absolute Error maintains the original scale of the data, making it potentially more interpretable in the action space’s context. MAE is also less sensitive to outlier predictions than squared error metrics, thus giving a more stable evaluation of typical performance. The maximum possible Brier score is 2, which is assigned to invalid predictions as a high penalty.
@@
  • Normalized Quantile Filtered Brier Mean Absolute Error This metric value is obtained by filtering out the Brier Absolute Errors that are lesser than or equal to the 5th percentile error or greater than or equal to the 95th percentile error, and then normalizing based on the quantile filtered minimum and maximum errors. This metric helps indicate how good or bad the majority of the predictions are. Discrepancy in trends between normalized and normalized quantile filtered Brier MAEs indicates the reliance on a few extremely good or extremely bad predictions.
@@
    This metric quantifies how much the worst-case error deviates from the typical (median) error produced by the model for a given subdataset of Procgen. A value significantly greater than 1 indicates the presence of some extremely bad predictions that can skew the results.
@@
    For our case, where it is a multi-class, single correct prediction task and the classes (actions) are mutually exclusive, false positives are equal to false negatives. Simply, an incorrectly predicted action is a false positive for its class, while also a false negative for the ground truth action class. Due to this, precision = recall = f1 score = exact match rate when calculating micro metrics in our case. The micro precision/recall/f1 is calculated over all timesteps of the test split of a given subdataset. This metric measures how precise the models are - high precision indicates that the model gets a high number of predictions correct, and a low precision suggests the model is frequently predicting actions that don’t match the ground truth.
  • Percentage invalids Predictions that fall outside the valid action space of a given subdataset of Procgen are considered invalid. In the case of the VLMs - GPT 4o, and GPT 4.1, the outputs that are not in the desired format of a vector the size of the action space with discrete action keys and their probabilities summing to 1 as its elements, are also considered invalid. Thus, percentage invalids is calculated as
@@
  • Micro Precision without Invalids This metric is a variation of micro precision where the false positives are only counted for valid predictions. By not considering the invalid predictions, this metric provides more insight into the model’s performance in the cases where it is able to produce valid outputs. A discrepancy in the trends between micro precision and micro precision without invalids indicates that a model would be more capable if it were fine-tuned, prompted, or had its output processed to constrain the predictions to the valid action space.
@@
    These metrics help to understand model biases and preferences towards specific action classes.
@@
    Macro $\text{Macro Precision}=\frac{1}{K}\sum_{k=1}^{K}\frac{TP_{k}}{TP_{k}+FP_{k}}$ (8)
    Macro $\text{Macro Recall}=\frac{1}{K}\sum_{k=1}^{K}\frac{TP_{k}}{TP_{k}+FN_{k}}$ (9)
[…]
    K here refers to the number of classes for a given subdataset. Macro precision, recall, and F1 metrics are calculated by averaging the class-wise metrics across all classes in the valid action space of a given subdataset for a given model. These metrics are crucial, as they help alleviate the effects of the dominant majority class that skews performance. With micro metrics, the majority classes are favored due to global aggregation. This leads to a biased view of performance, especially in the Procgen expert data, as there are numerous cases of class imbalance, which can skew the values of the micro metrics. Macro metrics prove to be a better indicator of a model’s performance when compared to micro metrics, irrespective of the nature of the test dataset, as they treat all classes equally regardless of size. Within this, Macro recall is typically the most representative metric for performance as it is neither affected by the majority classes, as in the case of micro metrics, nor is it affected by the performance of models on the rare classes, as in the case of macro precision or F1 scores.